A.) How many grams of CA2Cl2 must be used to provide 0.252 moles of chloride ions?

B.) Calculate the mass percent of C,H, and O in 1.205 moles of Vitamin C (C6H8O6).
Answer #1

Ans :

Each mol of Ca2Cl2 has two mole of chloride ions

So 0.252 moles of chloride ions will be formed from : 0.252 mol / 2 = 0.126 mol CaCl2

Molar mass of CaCl2 = 110.98 g/mol

So mass of CaCl2 to be used = mol x molar mass

= 0.126 mol x 110.98 g/mol

= 14.0 grams

B)

Molar mass of C6H8O6 = 176.12 g/mol

mass of C = 6 x 12.01 = 72.06 g

mass of H = 8 x 1.01 = 8.06 g

mass of O = 6 x 16 = 96 g

mass percent = ( mass of element in compound / molar mass) x 100

Mass % C = ( 72.06 / 176.12) x 100

= 40.92 %

Mass % H = (8.06 / 176.12) x 100

= 4.58 %

Mass % O = ( 96/176.12) x 100

= 54.51 %